# thanks: ORCID $0$-0001-9980-0363thanks: ORCID $0$-0001-9896-1018thanks: ORCID $0$-0003-1168-7883

### v1

@@ Abstract
The phenomenon of Heider (structural) balance is known for a long time (P. Bonacich and P. Lu, Introduction to Mathematical Sociology, Princeton UP, 2012). Yet it attracts attention of numerous computational scholars, as it is an example of a macroscopic ordering which emerges as a consequence of local interactions. In this paper, we investigate the thermal evolution (driven by thermal noise level $T$) of the work function $U(T)$ for Heider balance on several Archimedean lattices that contain separated triangles, pairs of triangles, chains of triangles and complex structures of triangles. To that end, the heat-bath algorithm is applied. Two schemes of link values updating are considered: synchronous and asynchronous. In the latter case, the analytical formula $U(T)=-\tanh(1/T)$ based on the partition function is provided. The Archimedean lattices are encoded with adjacency matrices, and Fortran procedures for their construction are provided. Finally, we present the mathematical proof that for any two-dimensional lattice, perfect structural (Heider) balance is unreachable at $T>0$.
@@ I Introduction
The sociophysical problem of structural balance [1, 2, 3, 4, 5, 6, 7]. governed by Heider’s algebra [1]:
@@
brought attention of physicists (see Refs. 8, 9 for recent reviews) as a three body problem with Hamiltonian
$\mathcal{H}=-\sum_{i}\sum_{j>i}\sum_{k>j}a_{ij}x_{ij}\cdot a_{jk}x_{jk}\cdot a_{ki}x_{ki},$ (1)
where $x_{ij}=\pm 1$ represent friendly and hostile relations between actors $i$ and $j$ and $a_{ij}\in\{0,1\}$ define the network/lattice topology.
Figure 1: Heider’s triads corresponding to balanced (1(a) and 1(c)) and imbalanced (1(b) and 1(d)) states
When in every triangle these relations obey Heider’s rules mentioned above, the system is termed as balanced in Heider’s sense—which means that only triangles presented in Figures 1(a) and 1(c) are observed in the system.
The earlier studies have been devoted to triangular [10], diluted [11] and densified triangulations [12]; chains of nodes [13]; classical random graphs [14]; and complete graphs [15, 16].
Recently, reaching structural balance has been investigated with thermal noise modeled either with heat-bath [17, 15, 16] or Glauber dynamics [18, 19]. A more refined dynamics based on the indirect reciprocity mechanism on a complete graph has been studied in Reference [20].
In this paper, we continue studies of the influence of thermal noise on the possibility of reaching structural balance on Archimedean lattices containing triangles (see Figure 2). There are eleven Archimedean lattices [21], which contain only regular polygons allowing for covering plane with tiles and keeping the translational symmetry of the lattice. The systematic names of these lattices come from the number of edges in tiles attached to every lattice node sorted in the possibly lowest lexicographic order—for example, we call lattice presented in Figure 2(b) $(3^{2},4,3,4)$ and not $(4,3^{2},4,3)$ or $(4,3,4,3^{2})$. In this way, the square lattice is called $(4^{4})$, the triangular lattice is called $(3^{6})$, the honeycomb lattice is called $(6^{3})$ and kagomé lattice is called $(3,6,3,6)$.
@@
Additionally, we inspect the system evolution for triangular $(3^{6})$ lattice (Figure 2(a)), diluted triangulations, and classical random graphs with mean node degree equal to $k=4$ and $k=5$, which also characterize the above-mentioned lattices. In Table 1 characteristics of the Archimedean lattices discussed in this work are collected. The social noise is modeled by the temperature parameter $T$ in the heat-bath scheme.
Our calculations are performed both with asynchronous and synchronous scheme. The former is known to reproduce properly the thermodynamic properties of systems in thermal equilibrium. The latter is justified if we intend to mimic the dynamics of opinions gathered periodically in polls. In this case, a respondent is not aware of the answers of other respondents during the same round. Such procedure of data collection has been applied in [22], to give an example.
The paper is organized as follows: in Section II we describe the methodology of studies (including the creation of adjacency matrices for the Archimedean lattices mentioned above, the heat-bath algorithm and calculation of the work function), Section III contains results of Monte Carlo simulations and Section IV is devoted to discussion of the obtained results and presentation of mathematical proof that structurally ordered phase is limited to $T=0$ on any two-dimensional lattice.
The paper is finalized with Appendix A showing lattices and adjacency matrices shapes and Appendix B with Fortran95 procedures for creating adjacency matrices for the lattices considered here.
@@ II Methods
We keep the simulation methodology as presented for diluted/densified triangulations and classical random graphs [11, 12, 14]. In this approach (1), every link $x_{ij}$ represents friendly ($x_{ij}=+1$) or hostile ($x_{ij}=-1$) attitudes between actors $i$ and $j$, while (binary and symmetric) adjacency matrix $\mathbf{A}=[a_{ij}]$ elements
@@
The Heider balance can be easily identified by checking the system work function [23, 24] per number of triangles $\Delta$ on lattice
$U/\Delta\equiv-\frac{\operatorname{\mathrm{Tr}}[(\mathbf{A}\circ\mathbf{X})^{3}]}{\operatorname{\mathrm{Tr}}(\mathbf{A}^{3})},$ (3)
where $\circ$ stands for the Hadamard product of the matrices and the matrix $\mathbf{X}=[x_{ij}]$. The system work function per triangle $U/\Delta$ is equal to $-1$ if and only if all triangles in the system are balanced.
II.1 Adjacency matrix representation of Archimedean lattices
In Appendix A in Figures 13(a), 10(a), 11(a), 12(a) and 9(a) the equivalents of the lattice shapes presented in Figure 2 are mapped onto distorted triangular lattices embedded in square lattices. The lattice edges are printed in thick solid black lines, while the underlying triangular lattice is presented in thin solid blue lines. The ‘unit cell’ is marked with a green dashed line. The periodic boundary conditions are assumed, and replicas of ‘elementary cell’ are marked with dotted cyan lines.
The Fortran procedures providing $\mathbf{A}$ for considered lattices are presented in Appendix B. Independently of the considered lattice, the module ‘par’ (see LABEL:lst:module-par in Appendix B) is required to define the linear size of the system $W$ and the lattice connectivity $k$.
Table 1: Characteristics of some Archimedean lattices. The second column indicates the lattice connectivity $k$. References to figures showing lattices’ shapes (and their shapes embedded in a distorted triangular lattice) are given in the third and the fourth columns. Figures showing the shapes of the adjacency matrix $\mathbf{A}$ are indicated in the fifth column. In the sixth column the labels of listings supplying $\mathbf{A}$ are listed
lattice $k$ shape shape in Appendix A matrix $\mathbf{A}$ in Appendix A subroutine in Appendix B
$(3^{6})$ 6 Figure 2(a) Figure 9(a) Figure 9(b) LABEL:lst:code-triangular
$(3^{2},4,3,4)$ 5 Figure 2(b) Figure 10(a) Figure 10(b) LABEL:lst:code-tri2al33434
$(3^{3},4^{2})$ 5 Figure 2(c) Figure 11(a) Figure 11(b) LABEL:lst:code-al33344
$(3^{4},6)$ 5 Figure 2(d) Figure 12(a) Figure 12(b) LABEL:lst:code-tri2mapleleaf
$(3,6,3,6)$ 4 Figure 2(e) Figure 13(a) Figure 13(b) LABEL:lst:code-tri2kagome
II.2 Heat-bath
The social noise level is mimicked by the temperature $T$ introduced in the probabilities of setting a hostile or friendly relation between actors. If $a_{ij}\neq 0$ then evolution of the link between nodes $i$ and $j$ of value $x_{ij}$ is given by
$x_{ij}(t+1)=\begin{cases}+1&\text{ with probability }p_{ij}(t),\\ -1&\text{ with probability }[1-p_{ij}(t)],\end{cases}$ (4a)
@@ II.2 Heat-bath
$p_{ij}(t)=\frac{\exp[\beta\xi_{ij}(t)]}{\exp[\beta\xi_{ij}(t)]+\exp[-\beta\xi_{ij}(t)]},$ (4b)
$\beta=1/T$ (4c)
and
$\xi_{ij}(t)=\sum_{k}a_{ik}x_{ik}(t)\cdot a_{kj}x_{kj}(t).$ (4d)
II.3 Calculations of $U(T)$
@@
$p(A\rightarrow B)=\frac{e^{-\beta E_{B}}}{\sum_{C}e^{-\beta E_{C}}}$ (5)
@@
$Z=\sum_{C}e^{-\beta E_{C}}.$ (6)
For one triad, we have $2^{3}$ microscopic states $C$, four of them with energy $U=-1$ and four with energy $U=+1$; hence the partition function is
@@
For two triads with a common link, we have $2^{5}$ microscopic states. The energy spectrum contains eight states of energy equal $-2$, eight states of energy equal $+2$, and sixteen states of energy equal zero. The direct summation over these states gives the partition function $Z_{2}=Z_{1}^{2}$. Accordingly,
$U_{2}=2U_{1}.$ (9)
@@
$Z_{N}=\sum_{i}\sum_{c}\exp(-\beta E(i,c))=2^{N}[2\cosh(\beta)]^{N},$ (10)
@@
where $i$ and $c$ are related to the states of internal and closing links, respectively. As a consequence, the energy per one triangle is equal to $-\tanh(\beta)$ again.
III Results
[…]
Figure 4: $\langle U_{\text{av}}\rangle/\Delta$ for synchronous (4(a), 4(c), 4(e), 4(g)), and asynchronous (4(b), 4(d), 4(f), 4(h)) update scheme. Nodes degree (on average, for the diluted triangular lattice and classical random graphs) $k=4$ (4(c), 4(d)), $k=5$ (4(e), 4(f)), $k=6$ (4(g), 4(h))
The energy per triangle defined in Equation 3 is therefore calculated in the final 90% of the time steps and then averaged to find $U_{\text{av}}$, and then averaged again over all the simulations performed. The results, denoted as $\langle U_{\text{av}}\rangle/\Delta$ are presented in LABEL:{fig:U_vs_T} as a function of temperature $T$, together with the curve given by
$U/\Delta=-\tanh(1/T),$ (11)
corresponding to the theoretical results of Section II.3.
The two approaches used in our calculations, synchronous and asynchronous, produce results that differ significantly at lower $T$. Figure 4(a) shows that the synchronous updating leads to $\langle U_{\text{av}}(T)\rangle/\Delta$ dependencies which are unique to each of the discussed Archimedean lattices. In the low-temperature limit, we observe perfectly balanced states, $\langle U_{\text{av}}\rangle/\Delta=-1$, in the $(3,6,3,6)$, $(3^{3},4^{2})$ and $(3^{4},6)$ lattices, but the decrease of energy with decreasing temperature occurs along a different curve for each of those lattices, and in all cases the values of the work function per triangle remain significantly larger than $-\tanh(1/T)$. On the other hand, only a slight decrease in the work function is found for the triangular lattice $(3^{6})$ and somewhat larger, to $\langle U_{\text{av}}(T\to 0)\rangle/\Delta\approx-0.5$, for the $(3^{2},4,3,4)$ lattice. For the asynchronous updating, the results in Figure 4(b) prove that for all types of the considered Archimedean lattices the temperature dependence of the work function per triad is exactly the same and identical to the theoretically predicted $U/\Delta=-\tanh(1/T)$.
For comparison, we match the Archimedean lattices with the corresponding diluted triangular lattices (DTL) [12] and classical random graphs (CRG) [14] having the same average node degree $k=4$, $5$, or $6$ as the connectivity of the considered Archimedean lattices. The only Archimedean lattice discussed here with $k=4$ is $(3,6,3,6)$ and for synchronous updating DTL and CRG of the same $k$ reveal similar but not identical characteristics, with non-perfect balance $\langle U_{\text{av}}(T)\rangle/\Delta\approx-0.9$ at $T\to 0$ as visible in Figure 4(c). It should be noted, though, that the random nature of the DTL and CRG graphs means that the simulations are performed on lattices with the same average node degree and number of nodes, but with different structure of connections (graph egdes) which results, among other things, in the number of triads in those lattices that is not constant in all simulations but found to be $\Delta=37\pm 3$ for DTL or $\Delta=11\pm 3$ for CRG, with $N=64$ in both cases. For $k=5$ (Figure 4(e)) and $k=6$ (Figure 4(g)) with synchronous updating the results for DTL and CRG are similar to those for $k=4$, which also means that they are not exactly in line with those for Archimedean lattices, in particular in the case of $(3^{2},4,3,4)$ lattice where the balanced state is not observed even in the low-temperature limit. At higher $T$, the results for all Archimedean, DTL and CRG lattices agree. For asynchronous updating, the results obtained for DTL and CRG at $k=4,5,6$ and presented in Figures 4(d), 4(f) and 4(h) do not differ from those of Archimedean lattices.
Figure 5: Values of $U_{\text{av}}/\Delta$ obtained in 100 simulations for each temperature $T$, for various simulation times $t_{\text{max}}=10^{4}$, $10^{6}$ and $10^{8}$, for (5(a)) $(3^{4},6)$ and (5(b)) $(3^{2},4,3,4)$ lattice and synchronous scheme of link updates
The dispersion of the values obtained in 100 simulations is usually relatively small, especially when asynchronous updating is applied. In most cases, it is also true for synchronous updating, except for the $(3^{2},4,3,4)$ and $(3^{4},6)$ lattices at $T<1$ where it was required to perform $t_{\text{max}}=10^{8}$ time steps, and for DTL and CRG lattices where $t_{\text{max}}=10^{6}$ was used. In particular in the case of $(3^{4},6)$ lattice shorter simulations do not allow us to see that the balanced state is achieved at $T\to 0$, as visible in Figure 5(a) showing the values $U_{\text{av}}/\Delta$ obtained in all performed simulations. Figure 5(b) confirms that on the $(3^{2},4,3,4)$ lattice the balanced state is not observed even in the low-temperature limit, and additionally the values of the work function obtained in the simulations performed are quantized and distributed over a relatively long interval. The values of the work function observed at the lowest $T$ are spaced in the intervals of $2/\Delta$, which corresponds to the difference related to the energetically nearest states. The same quantization of the work function applies for any other two-dimensional lattice that contains triangles as it is associated with changing a single term ($x_{ij}x_{jk}x_{ik}$—where $i$, $j$, $k$ are triangle vertices) in sum in nominator of Equation 3 from $+1$ to $-1$ (or vice versa).
Figure 6: Thermal evolution (in the low temperature limit $T<1$) of fractions of triads presented in Figure 1 in the case of asynchronous updating and different Archimedean lattices. The panels (6(a)), (6(b)), (6(c)), and (6(d)) correspond to the triangles presented in Figures 1(a), 1(b), 1(c), and 1(d), respectively. The error bars show the standard deviation of the results in 100 simulations. Dashed line for asynchronous updating (common result for all lattices)
If any of the systems discussed is in the balanced state, it means that all triads are balanced, that is, they are in one of two possible configurations: with all positive links ([$+$][$+$][ + + + ], Figure 1(a)) or with one positive and two negative links ([+][+--][ + - - ], Figure 1(c)). The fraction of triads of a given kind, $\langle f\rangle$, as a function of $T$ is shown in Figure 6. It reveals that in the cases where the balanced state is reached at $T\to 0$, about 25% of the triads are in the [$+$][$+$][ + + + ] configuration, and the rest are in the [+][+--][ + - - ] state. However, in that limit (only for $T<0.2$) those fractions vary substantially among the simulations, as shown by the error bars indicating the standard deviation, which is also the case for the $(3^{2},4,3,4)$ lattice which does not reach the balanced state and for all triads in that case including [++][++-][ + + - ] (Figure 1(b)) and [][---][ - - - ] (Figure 1(d)). On the other hand, the $(3^{6})$ triangular lattice produces almost identical results in all simulations. In the balanced state observed in the low-temperature limit, the fractions of triads obtained using the asynchronous updating scheme (dashed lines in Figure 6) are the same as for the balanced state reached with synchronous updating. However, in the case of asynchronous updating the increasing temperature causes much slower decrease of the number of balanced triads than for synchronous approach, which is also visible in slower increase of the work function with increasing $T$.
IV Discussion
@@
Here we show that the Heider model (1) on a two-dimensional lattice with the asynchronous updates is equivalent to a one-dimensional Ising model with the nearest-neighbor interactions.
@@
where $\sigma_{51}=\sigma_{1}$. This is a partition function of the one-dimensional nearest-neighbor Ising model with periodic boundary conditions. For $\sigma_{*}=-1$ we have

### v2

@@ Abstract
We investigate the work function $U(T)$ for the Heider balance, driven by a thermal noise $T$, on several planar networks that contain separated triangles, pairs of triangles, chains of triangles and complex structures of triangles. In simulations, the heat-bath algorithm is applied. Two schemes of link values updating are considered: synchronous and asynchronous (sequential). The latter results are compared with analytical calculations for small cliques. We argue that the actual shape of $U(T)$ is a consequence of a local topology rather than of a macroscopic ordering. Finally, we present the mathematical proof that for any planar lattice, perfect structural (Heider) balance is unreachable at $T>0$.
@@ I Introduction
The phenomenon of Heider (structural) balance is known for more than 70 years [1, 2]. It attracts continuous attention of numerous computational scholars, as it seems to be an example of a macroscopic ordering which emerges as a consequence of local interactions. From a sociological perspective, the problem deals with a cognitive dissonance, which appears when the partition of our acquaintances into enemies and friends remains fuzzy. From a computational point of view, the problem is to evaluate the stationary departure from a balanced state in the presence of noise. Such an algorithm could be used to predict conflicts, and perhaps to prevent them.
The sociophysical problem of structural balance [1, 3, 4, 5, 6, 7, 2] governed by Heider’s algebra [1]:
@@
attracted attention of physicists (see References 8, 9 for recent reviews) as a three body problem based on interaction in triads of actors.
Figure 1: Heider’s triads corresponding to balanced (1(a) and 1(c)) and imbalanced (1(b) and 1(d)) states. Solid blue lines and dashed red lines represent friendly ($+1$) and hostile ($-1$) relations, respectively
When relations in every triangle obey Heider’s rules mentioned above, the system is termed as balanced in Heider’s sense—which means that only triangles presented in Figures 1(a) and 1(c) are observed in the system. Measures of structural balance based on the concepts of weak and strong balance with application to several real-world signed networks are discussed in Reference 10, and further statistical perspective used to assess the empirical unbalanced patterns is provided in Reference 11 in particular for exponential random graphs.
The earlier studies have been devoted to triangular [12], diluted [13] and densified triangulations [14]; chains of nodes [15]; classical random graphs [16]; and complete graphs [17, 18]. Several works deal with a balance-imbalance phase transition at finite temperature [18, 19, 20, 21, 14, 22, 23]. In these works, numerical simulations are often backed by the mean-field approach [18, 20, 22]. The balance-imbalance phase transition is found to be of first order (e.g. in Erdős–Rényi networks [24]). In Reference 19, the authors have indicated that as the usual form of the work function consists of a sum over particular triads, the triads are independent even if they share a link. Below we analyze how this assertion is influenced by a local topology of the network.
Recently, reaching structural balance has been investigated with thermal noise modeled either with heat-bath [25, 17, 18] or Glauber dynamics [26, 27]. (A more refined dynamics based on the indirect reciprocity mechanism on a complete graph has been studied in Reference 28). In this paper, we continue studies of the influence of thermal noise on the possibility of reaching structural balance on Archimedean lattices, all of them planar (see Figure 2). The results of the simulations are accompanied by analytical calculations. The latter are also performed for cliques of four and five nodes.
Properties of triads of actors is a classic chapter of sociology since times of Georg Simmel [29]. The Archimedean lattices, as being periodic structures composed of triads, are convenient as a medium for computational research. There are eleven Archimedean lattices [30], which contain only regular polygons allowing for covering plane with tiles and keeping the translational symmetry of the lattice. The systematic names of these lattices come from the number of edges in tiles attached to every lattice node sorted in the possibly lowest lexicographic order—for example, we call lattice presented in Figure 2(b) $(3^{2},4,3,4)$ and not $(4,3^{2},4,3)$ or $(4,3,4,3^{2})$. In this way, the square lattice is called $(4^{4})$, the triangular lattice is called $(3^{6})$, the honeycomb lattice is called $(6^{3})$ and kagomé lattice is called $(3,6,3,6)$.
@@
Additionally, we inspect the system evolution for the triangular $(3^{6})$ lattice (Figure 2(a)), diluted triangulations, and classical random graphs with mean node degree equal to $k=4$ and $k=5$, which also characterize the above-mentioned lattices. The characteristics of the Archimedean lattices discussed in this work are collected in Table 1. Social noise is modeled by the temperature parameter $T$ in the heat-bath scheme.
To compare results of planar with non-planar lattices, we repeat the calculations for small cliques, i.e. fully connected networks, of $N$ nodes. In computational sociology, the structure of cliques is often used to explore properties of small groups [31, 32]. The case of $N=3$ (a triad) belongs to the set of Archimedean lattices. We will show that the obtained $U(T)$ dependence is the same for separate triads and for all Archimedean networks, where $U(T)$ is the work function defined in Section III.1. The cases of $N=4$ and $N=5$ are not planar. Yet, if a network is composed of cliques sharing only corners of the cliques, the obtained $U(T)$ is just a sum of work functions of individual cliques.
Our calculations for the cliques are analytical, and related to thermal equilibrium. The calculations for planar lattices are performed both with asynchronous and synchronous scheme. The former is known to reproduce properly the thermodynamic properties of systems in thermal equilibrium. The latter is sociologically justified if we intend to mimic the dynamics of opinions gathered periodically in polls. In this case, a respondent is not aware of the answers of other respondents during the same round. Such procedure of data collection has been applied in Reference 33, to give an example. Synchronous simulations allow us to identify metastable minima of the work function, which are visible as a departure from the asynchronous case.
The paper is organized as follows: in Section II we describe the methodology of the studies (including the creation of adjacency matrices for the Archimedean lattices mentioned above, the heat-bath algorithm and calculation of the work function), Section III contains results of theoretical calculations, presentation of the mathematical proof that structurally ordered phase is limited to $T=0$ on any planar lattice and Monte Carlo simulations, and Section IV is devoted to discussion of the obtained results.
The paper is finalized with Appendix A showing lattices and adjacency matrices shapes. Fortran95 procedures for creating adjacency matrices $\mathbf{A}$ for the lattices considered here are provided in the Supplemental Material [34].
@@ II Methods
We keep the simulation methodology as presented for diluted/densified triangulations and classical random graphs [13, 14, 16]. In this approach, the energy of the system is given by
$E=-\sum_{i}\sum_{j>i}\sum_{k>j}a_{ij}x_{ij}\cdot a_{jk}x_{jk}\cdot a_{ki}x_{ki},$ (1)
where every link $x_{ij}$ represents friendly ($x_{ij}=+1$) or hostile ($x_{ij}=-1$) attitudes between actors $i$ and $j$, while (binary and symmetric) adjacency matrix $\mathbf{A}=[a_{ij}]$ elements,
@@
The equivalents of the lattice shapes shown in Figure 2 mapped onto distorted triangular lattices embedded in square lattices and the corresponding adjacency matrices are presented in Appendix A, Figures 14, 11, 12, 13 and 10.
The Heider balance can be easily identified by checking the system energy [35, 36] per number of triangles $\Delta$ on the lattice,
$E/\Delta\equiv-\frac{\operatorname{\mathrm{Tr}}[(\mathbf{A}\circ\mathbf{X})^{3}]}{\operatorname{\mathrm{Tr}}(\mathbf{A}^{3})},$ (3)
where $\circ$ stands for the Hadamard product of the matrices and the matrix $\mathbf{X}=[x_{ij}]$. The system energy per triangle $E/\Delta$ is equal to $-1$ if and only if all triangles in the system are balanced.
Table 1: Characteristics of some Archimedean lattices. The second column indicates the lattice connectivity $k$. References to figures showing lattices’ shapes (and their shapes embedded in a distorted triangular lattice) are given in the third and the fourth columns. Figures showing the shapes of the adjacency matrix $\mathbf{A}$ are indicated in the fifth column. In the sixth column the labels of listings supplying $\mathbf{A}$ are listed. The latter are available in the Supplemental Material [34]
lattice $k$ shape shape in Appendix A matrix $\mathbf{A}$ in Appendix A subroutine in [34]
$(3^{6})$ 6 Figure 2(a) Figure 10(a) Figure 10(b) Listing 2
$(3^{2},4,3,4)$ 5 Figure 2(b) Figure 11(a) Figure 11(b) Listing 3
$(3^{3},4^{2})$ 5 Figure 2(c) Figure 12(a) Figure 12(b) Listing 4
$(3^{4},6)$ 5 Figure 2(d) Figure 13(a) Figure 13(b) Listing 5
$(3,6,3,6)$ 4 Figure 2(e) Figure 14(a) Figure 14(b) Listing 6
The social noise level is mimicked by the temperature $T$ introduced in the probabilities of setting a hostile or friendly relation between actors. If $a_{ij}\neq 0$ then evolution of the link between nodes $i$ and $j$ of value $x_{ij}$ is given by the heat-bath algorithm,
$x_{ij}(t+1)=\begin{cases}+1&\text{ with probability }p_{ij}(t),\\ -1&\text{ with probability }[1-p_{ij}(t)],\end{cases}$ (4)
@@
$p_{ij}(t)=\frac{\exp[\beta\xi_{ij}(t)]}{\exp[\beta\xi_{ij}(t)]+\exp[-\beta\xi_{ij}(t)]}$ (5)
with $\beta=1/T$, and
$\xi_{ij}(t)=\sum_{k}a_{ik}x_{ik}(t)\cdot a_{kj}x_{kj}(t).$ (6)
@@
$p(A\rightarrow B)=\frac{e^{-\beta(E_{B}-E_{A})}}{\sum_{C}e^{-\beta E_{C}}}$ (7)
@@
$Z=\sum_{C}e^{-\beta E_{C}},$ (8)
where $C$ denotes all microscopic states.
Knowing the partition function $Z$ we can determine the mean value of energy of the system as
$U=-\frac{d}{d\beta}\ln Z.$ (9)
III Results
III.1 Calculations
For one triad, we have $2^{3}$ microscopic states $C$, four of them with energy $E=-1$ and four with energy $E=+1$; hence the partition function is
@@ III.1 Calculations
For two triads with a common link, we have $2^{5}$ microscopic states. The energy spectrum contains eight states of energy equal $-2$, eight states of energy equal $+2$, and sixteen states of energy equal zero. The direct summation over these states gives the partition function $Z_{2}\propto Z_{1}^{2}$. Accordingly,
$U_{2}=2U_{1}$ (12)
and
$U_{2}/\triangle=-\tanh(\beta).$ (13)
@@
$Z_{\Delta}=\sum_{i}\sum_{c}\exp(-\beta E(i,c))=2^{\Delta+1}[2\cosh(\beta)]^{\Delta},$ (14)
@@
where $\Delta$ is the number of triads, while $i$ and $c$ are related to the states of internal and closing links, respectively. As a consequence, the mean energy per one triangle is equal to
$U_{\text{chain}}/\triangle=-\tanh(\beta)$ (15)
again.
Now let us consider a network of tetrahedrons, formed in a way that neighboring tetrahedrons have only common corners. An example of this structure in crystallography is YMn2 [37]. It is clear that states of links of different tetrahedra do not depend on each other. Therefore the system is equivalent to a set of separate tetrahedra. A direct enumeration of $642^{6}=$ 64 states of one tetrahedron gives the partition function
$Z_{\text{tetrahedra}}=8e^{-4\beta}+48+8e^{4\beta},$ (16)
which gives the expression per one triad
$U_{\text{tetrahedra}}(\beta)/\triangle=-\frac{e^{4\beta}-e^{-4\beta}}{e^{4\beta}+6+e^{-4\beta}}.$ (17)
With some patience, the same procedure can be applied for networks of cliques of five or more nodes. For five nodes in a clique we get the work function per one triad
$\begin{split}U_{\text{clique,5}}(\beta)/\triangle=\\ -\frac{z^{10}+4z^{4}+3z^{2}-3z^{-2}-4z^{-4}-z^{-10}}{z^{10}+10z^{4}+15z^{2}+12+15z^{-2}+10z^{-4}+z^{-10}},\end{split}$ (18)
where $z=e^{\beta}$. A comparison of functions $U(T)$ calculated from Equations 15, 17 and 18 is shown in Figure 4.
@@
The calculated dependencies of $U(T)$ (per triangle) for a single triangle (11), pairs of triangles (13), and the chain of triangles (15) are the same as for a one-dimensional Ising model with the nearest-neighbor interactions. As we show below, this is not only a coincidence, but the Heider model (1) on a two-dimensional lattice with the asynchronous updates is equivalent to a one-dimensional Ising model with the nearest-neighbor interactions [38, 39].
Figure 4: The plots $U(T)$ for the chain of triangles, for cliques of four nodes (tetrahedra) and for cliques of five nodes, obtained from Equations 15, 17 and 18
@@
where $\sigma_{51}=\sigma_{1}$. This is a partition function of the one-dimensional nearest-neighbor Ising model with periodic boundary conditions. For $\sigma_{*}=-1$, we have